# Learning Operators through Coefficient Mappings in Fixed Basis Spaces

### v1

@@ 3.2 Coefficient representations and learnability
Each row of $A$ corresponds to a projected training function in the coefficient space of $V_{\mathrm{in}}$.
A useful diagnostic for assessing the learnability of such coefficient representations is the effective rank of the matrix $A$.
@@
To gain further intuition, we conduct a simple experiment to illustrate the effect of the SVD cut-off on the effective rank of the coefficient matrix $A$ in the RFM setting. Specifically, we consider the one-dimensional Darcy flow problem introduced in [27]. For this test, we set the number of random features in both the input and output basis systems to $m_{1}=m_{2}=128$, and discretize the domain with $1000$ uniformly spaced grid points. Figure 9 shows the resulting coefficient distributions under different cut-off thresholds. We observe that as the cut-off increases, the coefficients become more concentrated, the variance across different modes decreases, and the effective rank correspondingly grows. This confirms that in the RFM setting, a larger cut-off can paradoxically lead to coefficient matrices with higher effective rank, owing to the rapid decay of the singular value spectrum.
Figure 6: *
Figure 7: *
Figure 8: *
Figure 9: Coefficient distributions. Illustration of the coefficient distributions obtained under different SVD cut-off thresholds in the RFM setting. As the cut-off increases, the coefficients become more concentrated, the variance across modes decreases, and the effective rank correspondingly grows.
As shown in Figure 10, although the training curves are nearly identical across different cut-off thresholds, the generalization performance exhibits a clear dependence on the effective rank of the coefficient matrix. When the cut-off is small, the resulting coefficients yield a low effective rank, indicating that the representation is concentrated in a limited number of directions. Such low-rank structure restricts expressiveness and leads to poor generalization. By contrast, larger cut-off thresholds increase the effective rank, producing coefficient representations that span a richer subspace of the function space. This higher effective rank is consistently associated with lower test error, demonstrating that effective rank is a key indicator of generalization performance in FB-C2CNet.
@@ 3.3 Output-Space-Induced Error Analysis
$NN_{\theta}:\bm{a}\in\mathbb{R}^{m_{1}}\to\bm{b}\in\mathbb{R}^{m_{2}},$
@@
Therefore, under the assumption that the FB-C2CNet is sufficiently well trained such that the learnable mapping $NN_{\theta}$ accurately captures the underlying operator $G$, the total approximation error is dominated by, and asymptotically bounded by, the projection error associated with the output space $V_{\mathrm{out}}$, i.e.,
@@ 4 Numerical Experiments
In this section, we present a series of experiments across a broad spectrum of PDE settings to demonstrate the effectiveness of our method: (i) the Darcy flow problem on regular domains in both one and two dimensions, (ii) the Poisson equation on a two-dimensional complex domain, (iii) 2D Darcy flow on a complex domain, (iv) elastic plate problems, and (v) high-dimensional Poisson equations on regular domains. These examples encompass diverse dimensionalities, domain geometries, and operator learning scenarios, following the setups in [35, 36].
@@
Unless otherwise stated, the neural operator is implemented as a fully connected network with architecture $m_{1},512,m_{2}]$, where $m_{1}$ represents the number of basis functions used to approximate the input function space, and $m_{2}$ denotes the number of basis functions used to approximate the output function space. All models are trained with the Adam optimizer, employing a learning rate schedule that decays from $10^{-3}$ to $10^{-6}$ (as illustrated in Figure 58). All experiments are performed on a single NVIDIA RTX 4090 GPU. We utilize the Gmsh Python library for mesh generation. The resulting mesh is then used to construct the FEM basis functions. Additional implementation details and hyperparameter settings are provided in the Appendix.
@@
$\frac{d}{dx}\left(a(u)\frac{du}{dx}\right)$ $=f(x),\quad x\in[0,1],$ (27)
@@
where the solution-dependent permeability is $a(u(x))=0.2+u^{2}(x)$ and the input term $f$ is modeled as a Gaussian random field, $f\sim\mathcal{GP}(0,k(x,x^{\prime}))$, with covariance kernel $k(x,x^{\prime})=\sigma^{2}\exp\left(-\frac{|x-x^{\prime}|^{2}}{\ell^{2}}\right)$, $\ell=0.04$, and $\sigma=1.0$. The dataset is from [36, 37]. Our objective is to learn the map:
@@
The training set consists of 500 samples and the testing set consists of 200 samples. The spatial domain $[0,1]$ is discretized uniformly with 2000 grid points.
C2C vs. P2C: Efficiency and Accuracy. In this example, we compare the coefficient-to-coefficient (C2C) and point-to-coefficient (P2C) approaches using basis functions from the finite element method and the random feature model. For the P2C setting, the network architecture and training procedure remain identical to those of the C2C framework, except that the neural network takes pointwise function values as inputs instead of coefficients. This design highlights the effectiveness of using coefficient representations as neural network inputs. As shown in Figure 13, with 2000 spatial points, the P2C method requires a high-dimensional input (dimension 2000), resulting in a larger number of network parameters and more difficult training, which leads to slower loss decay and lower accuracy. In contrast, the proposed C2C method uses far fewer parameters and achieves higher accuracy, highlighting its efficiency.
Figure 11: *
@@
Figure 13: 1D Darcy Flow: Training curve comparison between the C2C and P2C methods with different basis functions. The number of grid points is set to $n_{1}=n_{2}=2000$, and the number of basis functions is $m_{1}=128$ and $m_{2}=32$.
Generalization through different resolutions. We train both RFM-C2C and FEM-C2C on data generated with $n_{\mathrm{train}}=100$ grid points, and then evaluate their performance on test datasets with $n=40$, $500$, $1000$, and $2000$ grid points, respectively. This experiment examines the ability of the learned operators to generalize across different spatial resolutions. The results, summarized in Fig. 14, show that the proposed RFM-C2C maintains stable accuracy across unseen resolutions.
Figure 14: 1D Darcy Flow: FB-C2CNet is trained on a uniform $100$-point grid and tested on different resolutions ($n=40,500,1000,2000$). The network exhibits strong resolution invariance with consistent test accuracy across all cases.
4.2 2D Darcy Flow in regular domain
In this example, we consider a two-dimension Darcy flow equation in the square domain $\Omega=[0,1]^{2}$. It is defined as
@@
The diffusion coefficient $a(x,y)$ is sampled from random field $exp(\psi)$, where $\psi=\mathcal{N}(0,(-\Delta+9I)^{-2})$. The dataset is from [9, 35]. The training set consists of 2400 samples and the testing set consists of 600 samples. The spatial domain $0,1]\times[0,1]$ is discretized uniformly with $141\times 141$ grid points.
C2C vs. P2C: Efficiency and Accuracy. In this example, similar to the 1D Darcy flow case, we compare the C2C and P2C approaches with different choices of basis functions. As shown in Figure 17, the results indicate that the C2C method achieves lower training error, which in turn leads to higher accuracy.
Figure 15: *
@@
Figure 17: 1D Darcy Flow: Training curve comparison between the C2C and P2C methods with different basis functions. The number of grid points is set to $n_{1}=n_{2}=141\times 141$, and the number of basis functions is $m_{1}=1024$ and $m_{2}=2048$.
4.3 Poisson equation in complex domain
In this example, we consider a two-dimensional Poisson equation defined as
@@ 4.3 Poisson equation in complex domain
where the computational domain $\Omega$ is a rectangle with three interior holes of different shapes. The objective is to learn the solution operator
@@
In this experiment, we employ both FEM-C2C and RFM-C2C. The finite element results are shown in Figure 20 and Figure 24, which illustrate the worst-case comparison in the test dataset together with the corresponding pointwise error. Similarly, the RFM results are presented in Figure 27 and Figure 31, also demonstrating the worst-case comparison in the test dataset and the associated pointwise error.
Relation between Basis Approximation and FB-C2C Error. In this experiment, we aim to verify that the accuracy of the proposed FB-C2C method is influenced by the approximation capability of the output basis system. As illustrated in Figure 20(b) and Figure 27(b), the red dashed line represents the approximation error of the output basis system to the output function space, while the blue solid line represents the accuracy of the FB-C2C method in solving the problem. It can be observed that the blue curve always lies above the red dashed curve, indicating that the accuracy of FB-C2C is bounded by the approximation accuracy of the output basis system (i.e., the error cannot be smaller than the approximation error). In this experiment, increasing the number of output basis functions $m_{2}$ generally improves the approximation capability of the output basis system. Consequently, as $m_{2}$ increases, both the approximation error and the FB-C2C error decrease. The fact that the two curves nearly coincide further demonstrates that FB-C2C is sufficiently trained in this setting.
Figure 18: *
(a)FEM basis for output function space.
Figure 19: *
(b) FEM basis.
Figure 20: Poisson equation in complex domain: FEM results. (a) Finite element mesh of the output basis system. (b) The FEM-C2C error (blue) is bounded below by the approximation error of the output basis system (red), and both decrease as the number of output bases $m_{2}$ increases.
Figure 21: *
Figure 22: *
Figure 23: *
Figure 24: Poisson equation in complex domain: Worst-case sample from the complex domain Poisson dataset solved by FEM-C2C. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
Figure 25: *
@@
Figure 26: *
@@
Figure 27: Poisson equation in complex domain: RFM results. (a) Distribution of sampled points. (b) The RFM-C2C error (blue) is bounded below by the approximation error of the output basis system (red), and both decrease as the number of output bases $m_{2}$ increases.
Figure 28: *
Figure 29: *
Figure 30: *
Figure 31: Poisson equation in complex domain: Worst-case sample from the complex domain Poisson dataset solved by RFM-C2C. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
@@ 4.4 Elastic plate
In this example, we demonstrate the capability of the proposed method to approximate solution operators for problems with coupled multi-field outputs. This investigation concerns a thin rectangular plate subjected to in-plane loading, described by the two-dimensional plane stress elasticity equations under the assumption of negligible body forces. The computational domain is defined as $\Omega=[0,1]\times[0,1]$ with a thin rectangular region removed from the interior, representing a cut-out in the plate. The governing equations are
$\begin{array}[]{cc}\nabla\cdot\mathbf{\sigma}+\mathbf{b}(\mathbf{x})=0,\quad\mathbf{x}=(x,y),&\\ (u,v)=0,\quad\forall\,x=0,&\\ \sigma\cdot\mathbf{n}=f(\mathbf{x}),\quad\forall\,x=1,\end{array}$ (30)
where $\sigma$ denotes the Cauchy stress tensor, $\mathbf{b}(\mathbf{x})=0$ is the body force, $\mathbf{n}$ is the unit outward normal vector, $f(\mathbf{x})$ is the boundary load, and $u(\mathbf{x})$ and $v(\mathbf{x})$ represent the displacements in the $x$- and $y$-directions, respectively. The material parameters are the Young’s modulus $E=300\times 10^{5}$ and Poisson’s ratio $\nu=0.3$. Under plane stress conditions, the constitutive relation between stress and displacement gradients is given by
$\begin{bmatrix}\sigma_{xx}\\ \sigma_{yy}\\ \tau_{xy}\end{bmatrix}=\frac{E}{1-\nu^{2}}\begin{bmatrix}1&\nu&0\\ \nu&1&0\\ 0&0&\frac{1-\nu}{2}\end{bmatrix}\begin{bmatrix}\displaystyle\frac{\partial u}{\partial x}\\[8.61108pt] \displaystyle\frac{\partial v}{\partial y}\\[8.61108pt] \displaystyle\frac{\partial u}{\partial y}+\frac{\partial v}{\partial x}\end{bmatrix}$. (31)
The boundary loading $f(1,y)$ applied to the right edge of the plate is modeled as a Gaussian random field. Our objective is to learn the map:
@@
The dataset for this problem is adopted from [36]. The input grid points are uniformly sampled over the domain with $101$ points, while the test grid contains $1048$ points, which is shown in Figure 35 (a). In total, $1850$ samples are used for training and $100$ samples are reserved for testing.
@@
Scalar-Valued vs. Vector-Valued Basis Functions: Efficiency and Accuracy. Both methods use $m_{1}=64$ input basis functions, a count precisely set in one dimension. Their output dimensions differ: in the 2D scalar-valued case, RFM attains $m_{2}=512$ while FEM yields $m_{2}=540$, as the latter cannot precisely mesh 512 triangles; in the vector-valued case, randomness generates a different basis, but these $m_{2}$ values are preserved. Two different meshes for the FEM basis generation are shown in Figure 35 (b) and Figure 35 (c).
As the vector-valued basis requires only half the output coefficients in this example, it contains significantly fewer trainable parameters, resulting in faster convergence and enhanced training efficiency. The relative error curves in Figure 38 (RFM) show similar accuracy for scalar-valued and vector-valued bases. Conversely, Figure 41 (FEM) reveals a significant accuracy gap, with the vector-valued basis performing much worse than the scalar-valued one. Therefore, the vector-valued basis provides superior training efficiency over the scalar-valued basis and, as realized in the RFM-C2C method, markedly better accuracy than the vector-valued FEM-C2C method.
Figure 32: *
(a) Sample grid points.
Figure 33: *
(b) $\psi$ $(\psi^{v})$ FEM basis mesh.
Figure 34: *
@@
Figure 38: Training curve comparison between the scalar-valued and vector-valued RFM basis. The number of grid points is set to $n_{1}=101$ and $n_{2}=1048$, and the number of basis functions is $m_{1}=64$ and $m_{2}=512$.
Figure 39: *
@@
Figure 41: Training curve comparison between the scalar-valued and vector-valued FEM basis. The number of grid points is set to $n_{1}=101$ and $n_{2}=1048$, and the number of basis functions is $m_{1}=64$ and $m_{2}=540$.
4.5 2D Darcy Flow in complex domain
We investigate Darcy flow in a two-dimensional L-shaped region to validate our approach in handling complex domain configurations. A distinctive feature of this example is that it involves a mapping from two functions to a single function, thereby highlighting the broad applicability of our method. The problem is formulated as

### v2

@@ 3.2 Coefficient representations and learnability
Each row of $A$ corresponds to a projected training function in the coefficient space of $V_{\mathrm{in}}$. A useful diagnostic for assessing the learnability of such coefficient representations is the effective rank of the matrix $A$.
@@
To gain further intuition, we conduct a simple experiment to illustrate the effect of the SVD cut-off on the effective rank of the coefficient matrix $A$ in the RFM setting. Specifically, we consider the one-dimensional Darcy flow problem introduced in [27] (Details can be found in Section 4.1). For this test, we set the number of random features in both the input and output basis systems to $m_{1}=m_{2}=128$, and discretize the domain with $1000$ uniformly spaced grid points. Figure 4 shows the resulting coefficient distributions under different cut-off thresholds. We observe that as the cut-off increases, the coefficients become more concentrated, the variance across different modes decreases, and the effective rank correspondingly grows. This confirms that in the RFM setting, a larger cut-off can paradoxically lead to coefficient matrices with higher effective rank, owing to the rapid decay of the singular value spectrum.
Figure 4: Coefficient distributions. Illustration of the geometric distributions of the coefficient matrix obtained under different SVD cut-off thresholds in the RFM setting. For each cut-off value, we plot the histogram of nearest-neighbor distances in the high-dimensional coefficient space, where each bar corresponds to the empirical distribution of Euclidean distances between one sample and its closest distinct neighbor among the $3200$ samples in $A\in\mathbb{R}^{3200\times 300}$. As the cut-off threshold increases, the coefficients become more concentrated, the variance across modes decreases, and the effective rank correspondingly grows.
As shown in Figure 5, although the training curves are nearly identical across different cut-off thresholds, the generalization performance exhibits a clear dependence on the effective rank of the coefficient matrix. When the cut-off is small, the resulting coefficients yield a low effective rank, indicating that the representation is concentrated in a limited number of directions. Such low-rank structure restricts expressiveness and leads to poor generalization. By contrast, increasing the cut-off raises the effective rank, producing coefficient representations that span a richer subspace of the function space. This increase in effective rank is consistently associated with improved generalization performance, demonstrating that effective rank is a key indicator of generalization in FB-C2CNet. However, overly large cut-off values are also undesirable, as they overly concentrate coefficients and discard important structural information in the data.
@@ 3.3 Output-Space-Induced Error Analysis
$NN_{\theta}:\bm{a}\in\mathbb{R}^{m_{1}}\mapsto\bm{b}\in\mathbb{R}^{m_{2}},$
@@
Therefore, under the assumption that the FB-C2CNet is sufficiently well trained such that the learnable mapping $NN_{\theta}$ accurately captures the underlying operator $G$, the total approximation error is dominated by and bounded by the projection error associated with the output space $V_{\mathrm{out}}$, i.e.,
@@ 4 Numerical Experiments
In this section, we present a series of experiments across a broad spectrum of PDE settings to demonstrate the efficiency and effectiveness of our method: (i) the Darcy flow problem on regular domains in both one and two dimensions, (ii) the Poisson equation on a two-dimensional complex domain, (iii) 2D Darcy flow on a complex domain, (iv) elastic plate problems, (v) time dependent nonlinear KdV–Burgers equation, (vi) Helmholtz equation with only one weak solution, (vii) high-dimensional Poisson equations on regular domains. Inverse boundary problem for the Stokes system will be presented in next section. These examples encompass diverse dimensionalities, domain geometries, and operator learning scenarios. Datasets for experiments (i-v) follow [24, 25], data setups for other examples are described separately in their respective subsections.
@@
Unless otherwise stated, the neural operator is implemented as a fully connected network with architecture $m_{1},-,m_{2}]$, where $m_{1}$ and $m_{2}$ are the input and output dimensions determined by the number of basis functions used to approximate the input function space and output function space, and – represents a cascade of hidden layers. The depth and width of these hidden layers are not fixed; instead, they are designed to maximize representational capacity under a parameter budget of 500,000. All models are trained with Adam optimizer, employing a learning rate schedule that decays from $10^{-2}$ to $10^{-4}$ (as illustrated in Figure 25). All experiments were performed on a single NVIDIA RTX 4090 GPU. We utilize the Gmsh Python library for mesh generation. The resulting mesh is then used to construct the FEM linear basis functions. Additional implementation details and hyperparameter settings are provided in the Appendix.
@@
$\frac{d}{dx}\left(a(u)\frac{du}{dx}\right)$ $=f(x),\quad x\in(0,1),$ (27)
@@
where the solution-dependent permeability is $a(u(x))=0.2+u^{2}(x)$ and the input term $f$ is modeled as a Gaussian random field, $f\sim\mathcal{GP}(0,k(x,x^{\prime}))$, with covariance kernel $k(x,x^{\prime})=\sigma^{2}\exp\left(-\frac{|x-x^{\prime}|^{2}}{\ell^{2}}\right)$, $\ell=0.04$, and $\sigma=1.0$. The dataset is from [25, 2]. The training set consists of 800 samples and the testing set consists of 200 samples. The spatial domain $[0,1]$ is discretized uniformly with 2000 grid points. Our objective is to learn the map:
@@
C2C vs. P2C: Efficiency and Accuracy. In this example, we compare the coefficient-to-coefficient (C2C) and point-to-coefficient (P2C) approaches using basis functions from the random feature model and the finite element method. For the P2C setting, the network architecture and training procedure remain identical to those of the C2C framework, except that the neural network takes pointwise function values as inputs instead of coefficients. It means that the input layer dimension of neural network in the P2C method is equal to the number of spatial grid points ($n_{1}=n_{2}=2000$ in this case), while the C2C method uses a significantly smaller input dimension determined by the number of basis functions ($m_{1}=m_{2}=128$ in this case). This design highlights the effectiveness of using coefficient representations as neural network inputs. As shown in Figure 6, with 2000 spatial points, the P2C method requires a high-dimensional input (dimension 2000), resulting in a larger number of network parameters and more difficult training, which leads to slower loss decay and lower accuracy. In contrast, the proposed C2C method uses far fewer parameters and achieves higher accuracy, highlighting its efficiency.
@@
Figure 6: 1D Darcy Flow: Training curve comparison between the C2C and P2C methods with different basis functions. The number of grid points is set to $n_{1}=n_{2}=2000$, and the number of basis functions is $m_{1}=128$ and $m_{2}=128$.
Generalization through different resolutions. From the original set of $n_{1}=n_{2}=2000$ points, we take every 20th point, resulting in a subset of 100 points used for training. We train both RFM-C2C and FEM-C2C on data with $n_{1}=n_{2}=100$ grid points, and then evaluate their performance on test datasets with $n_{1}=n_{2}=40$, $500$, $1000$, and $2000$ grid points, respectively. This experiment examines the ability of the learned operators to generalize across different spatial resolutions. The results, summarized in Figure 7, show that the proposed methods maintain stable accuracy across unseen resolutions and RFM-C2C achieves better performance.
Figure 7: 1D Darcy Flow: FB-C2CNet is trained on a uniform $100$-point grid and tested on different resolutions ($n_{1}=n_{2}=40,500,1000,2000$). The network exhibits strong resolution invariance with consistent test accuracy across all cases.
Table 2: 1D Darcy Flow: The comparison of the C2C framework and the P2C framework in two different RFM and FEM basis spaces. The cost time for the encoder means the time required to generate the basis functions and project the input functions to the corresponding basis space. The training time refers to the time required to train the neural operator.
Cost time Training time Test relative $L^{2}$ error Neural network structure
Basis Encoder C2C P2C C2C P2C C2C P2C
[…]
4.2 Darcy Flow in a rectangular domain
In this example, we consider a two-dimensional Darcy flow equation in the square domain $\Omega=(0,1)^{2}$. It is defined as
@@ 4.2 Darcy Flow in a rectangular domain
The diffusion coefficient $a(x,y)$ is sampled from random field $exp(\psi)$, where $\psi=\mathcal{N}(0,(-\Delta+9I)^{-2})$. The dataset is from [31, 24]. The training set consists of 2400 samples, and the testing set consists of 600 samples. The spatial domain $0,1]\times[0,1]$ is discretized uniformly with $141\times 141$ grid points.
C2C vs. P2C: Efficiency and Accuracy. In this example, similar to the 1D Darcy flow case, we compare the C2C and P2C approaches with different choices of basis functions. The learning rate for them starts from 1e31e-3 and is multiplied by 0.9 every 400 steps for a total of 10000 training steps. As shown in Figure 8 and Table 3, the results indicate that the C2C framework achieves lower training time since it has fewer parameters to be trained than the P2C framework, which in turn leads to higher accuracy in both basis functions. For the encoder part, generating FEM basis functions on a 2D grid is more computationally expensive than in 1D, making the efficiency advantage of the RFM basis even more pronounced in this case.
@@
Figure 8: Darcy Flow in a rectangular domain: Training curve comparison between the C2C and the P2C frameworks with different basis functions. The number of grid points is set to $n_{1}=n_{2}=141\times 141$. The numbers of basis functions are $(m_{1},m_{2})=(512,2048)$ for the RFM basis, and $(m_{1},m_{2})=(553,2059)$ for the FEM basis.
Table 3: Darcy Flow in a rectangular domain: The comparison of the C2C framework and the P2C framework in two different RFM and FEM basis spaces. The cost time for the encoder means the time required to generate the basis functions and project the input functions to the corresponding basis space. The training time refers to the time required to train the neural operator.
Cost time Training time Test relative $L^{2}$ error Neural network structure
Basis Encoder C2C P2C C2C P2C C2C P2C
RFM 3.67s 274.40s 302.11s 6.74e-3 39.90e-3 [512,180,180,2048] $141^{2}$,180,180,2048]
FEM 447.05s 291.17s 317.65s 6.92e-3 36.00e-3 [553,180,180,2059] $141^{2}$,180,180,2059]
Comparison with different operator learning methods. For a comprehensive evaluation, we benchmark our approach against four representative operator‑learning methods: BasisONet, B2B, DeepONet, and FNO2D. Their training times (on the same training set), network parameter counts, and mean relative $L^{2}$ errors ($RL^{2}E$) / mean squared errors ($MSE$) (on the same test set) are reported in Table 4. In comparison with other methods, the two FB-C2C methods train significantly faster, except against DeepONet. However, DeepONet exhibits substantially larger errors. Although FNO2D achieves slightly lower errors than our method, it requires far more parameters and longer training times. Additional comparative results on other benchmark problems are provided in the Appendix.
Table 4: Darcy Flow in a rectangular domain: Performance comparison for 2D Darcy flow in a rectangular domain. For C2C method, the total time is the sum of the encoder time and the training time.
Time Params $RL^{2}E$ MSEMSE
BasisONet [24] 88m1s 2,712,174 7.02e-3 9.12e-8
B2B [25] 94m2s 499,758 11.74e-3 27.31e-8
DeepONet [25] 4m10s 504,926 115.14e-3 2588.68e-8
FNO2D [34] 31m22s 1,188,353 5.43e-3 5.19e-8
RFM-C2C 5s+3m17s 492,864 6.47e-3 8.68e-8
FEM-C2C 7m27s+4m51s 504,979 6.92e-3 9.03e-8
4.3 Poisson equation in the complex domain
In this example, we consider a two-dimensional Poisson equation on the complex domain $\Omega$ defined as
@@ 4.3 Poisson equation in the complex domain
where the domain $\Omega$ is a rectangle with three interior holes of different shapes. The objective is to learn the solution operator
@@
In this experiment, we employ both FEM-C2C and RFM-C2C methods. The training set consists of 1600 samples and the testing set consists of 400 samples. The basis numbers for encoder and decoder are set to $(m_{1},m_{2})=(533,1586)$ and $(m_{1},m_{2})=(512,1600)$, respectively. We adopt a piecewise exponential learning rate decay: The learning rate for the FEM-C2C method starts from 1e51e-5 and is multiplied by 0.9 every 1000 steps for a total of 30000 training steps. The initial rates are 1e31e-3, 1e41e-4, and 1e51e-5 for three consecutive 10000-step segments, with a decay factor of 0.9 applied every 1000 steps within each segment for RFM-C2C. The FEM basis meshes for encoder and decoder are shown in Figure 9 (b) and (c). The mean relative $L^{2}$ errors on the test set is 4.13e34.13e-3 for FEM-C2C and 7.69e37.69e-3 for RFM-C2C method. The worst case is shown in Figure 10 and Figure 11, respectively.
Relation between Basis Approximation and FB-C2C Error. In this experiment, we aim to verify that the accuracy of the proposed FB-C2C method is influenced by the approximation capability of the output basis system. As illustrated in Figure 12 (a) and (b), the red dashed line represents the approximation error of the output basis system to the output function space, while the blue solid line represents the accuracy of the FB-C2C method in solving the problem. It can be observed that the blue curve always lies above the red dashed curve, indicating that the accuracy of FB-C2C is bounded by the approximation accuracy of the output basis system (i.e., the error cannot be smaller than the approximation error). In this experiment, increasing the number of output basis functions $m_{2}$ generally improves the approximation capability of the output basis system. This process is mathematically formulated in the previous inequality (26). Consequently, as $m_{2}$ increases, both the approximation error and the FB-C2C error decrease. The fact that the two curves nearly coincide further demonstrates that FB-C2C is sufficiently trained in this setting.
@@
(b) FEM mesh for encoder.
(c) FEM mesh for decoder.
Figure 9: 2D Poisson equation: Data representation and FEM mesh discretization. (a) Sample grid points for the functions. (b) FEM mesh for the encoder. (c) FEM mesh for the decoder.
Figure 10: 2D Poisson equation: Worst-case sample from the complex domain Poisson dataset solved by FEM-C2C. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
Figure 11: 2D Poisson equation: Worst-case sample from the complex domain Poisson dataset solved by RFM-C2C. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
(a)FEM basis.
@@
Figure 12: 2D Poisson equation: Prediction error of the learned operator versus the best possible projection error in the fixed basis space. The C2C error (blue) is bounded below by the approximation error of the output basis system (red), and both decrease as the number of output bases $m_{2}$ increases. (a) FEM basis. (b) RFM basis.
@@ 4.4 Elastic plate
In this example, we demonstrate the capability of the proposed method to approximate solution operators for problems with coupled multi-field outputs. This investigation concerns a thin rectangular plate subjected to in-plane loading, described by the two-dimensional plane stress elasticity equations under the assumption of negligible body forces. The computational domain is defined as $\Omega=(0,1)\times(0,1)$ with a thin rectangular region removed from the interior, representing a cut-out in the plate. The equilibrium equations are
$\begin{array}[]{cc}\nabla\cdot\mathbf{\sigma}+\mathbf{b}(\mathbf{x})=\mathbf{0},\quad\mathbf{x}=(x,y)\in\Omega,\end{array}$ (30)
where $\mathbf{\sigma}=\left[\begin{matrix}\sigma_{xx}&\tau_{xy}\\ \tau_{xy}&\sigma_{yy}\end{matrix}\right]$ denotes the Cauchy stress symmetric tensor ($\tau_{xy}=\tau_{yx}$), $\mathbf{b}(\mathbf{x})=\mathbf{0}$ is the body force. The material parameters are the Young’s modulus $E=300\times 10^{5}$ and Poisson’s ratio $\nu=0.3$. Under plane stress conditions, the constitutive relation between stress and displacement gradients is given by
$\begin{bmatrix}\sigma_{xx}\\ \sigma_{yy}\\ \tau_{xy}\end{bmatrix}=\frac{E}{1-\nu^{2}}\begin{bmatrix}1&\nu&0\\ \nu&1&0\\ 0&0&\frac{1-\nu}{2}\end{bmatrix}\begin{bmatrix}\epsilon_{xx}\\ \epsilon_{yy}\\ \gamma_{xy}\end{bmatrix}$. (31)
The kinematic equations relate the displacements to the strains as
$\epsilon_{xx}=\frac{\partial u}{\partial x},\quad\epsilon_{yy}=\frac{\partial v}{\partial y},\quad\gamma_{xy}=\frac{\partial u}{\partial y}+\frac{\partial v}{\partial x}.$ (32)
The boundary conditions consist of fixed displacements on the left edge of the plate and applied traction on the right edge, expressed as
[…]
where $\mathbf{n}$ is the unit outward normal vector on the boundary, $\mathbf{f}(\mathbf{x})=[0,f(1,y)]^{T}$ is the boundary load applied at the edge $x=1$, and $u(\mathbf{x})$ and $v(\mathbf{x})$ represent the displacements in the $x$- and $y$-directions, respectively. The boundary loading $f(1,y)$ applied to the right edge of the plate is modeled as a Gaussian random field. Our objective is to learn the map:
@@
The dataset for this problem is adopted from [25]. The input grid points are uniformly sampled over the domain with $101$ points, while the output grid contains $1048$ points, which is shown in Figure 13 (a). In total, $1850$ samples are used for training and $100$ samples are reserved for testing.
@@
Scalar-Valued vs. Vector-Valued Basis Functions: Efficiency and Accuracy. Both methods use $m_{1}=64$ input basis functions, a count precisely set in one dimension. Their output dimensions differ: in the 2D scalar-valued case, RFM attains $m_{2}=512$ while FEM yields $m_{2}=540$, as the latter cannot precisely mesh 512 triangles; in the vector-valued case, randomness generates a different basis, but these $m_{2}$ values are preserved. When we generate scalar valued basis functions, we just set $\psi_{i}=\psi_{i}^{u}$. Two different meshes for the FEM basis generation are shown in Figure 13 (b) and Figure 13 (c).
Some experimental results are shown in Table 5. As the vector-valued basis requires only half the output coefficients in this example, it contains significantly fewer trainable parameters, resulting in faster convergence and enhanced training efficiency. The relative error curves in Figure 14 (RFM) show similar accuracy for scalar-valued and vector-valued bases. Conversely, Figure 15 (FEM) reveals a significant accuracy gap, with the vector-valued basis performing much worse than the scalar-valued one. Therefore, the vector-valued basis provides superior training efficiency over the scalar-valued basis and, as realized in the RFM-C2C method, markedly better accuracy than the vector-valued FEM-C2C method.
Table 5: Elastic plate: The comparison of C2C framework and P2C framework in scalar-valued and vector-valued RFM and FEM basis spaces. The cost time for encoder means the time required to generate the basis functions and project input functions to the corresponding basis space. The training time refers to the time required to train the neural operator.
Cost time Training time Test relative $L^{2}$ error Neural network structure
Basis Encoder C2C P2C C2C P2C C2C P2C
Scalar RFM 0.82s 43.40s 48.11s 8.47e-3 141.56e-3 [64,512,1024] [101,512,1024]
Scalar FEM 11.75s 49.01s 52.80s 1.68e-3 1.93e-3 [64,512,1080] [101,512,1080]
Vector RFM 0.82s 42.11s 64.86s 7.81e-3 140.81e-3 [64,512,512] [101,512,512]
Vector FEM 11.75s 40.22s 42.84s 626.44e-3 626.43e-3 [64,512,540] [101,512,540]
(a) Sample grid points for output functions.
(b) $\psi^{v}$ FEM basis mesh.
@@
Figure 14: Elastic plate: Training curve comparison between the scalar-valued and vector-valued RFM basis. The number of grid points is set to $n_{1}=101$ and $n_{2}=1048$, and the number of basis functions is $m_{1}=64$ and $m_{2}=512$.
@@
Figure 15: Elastic plate: Training curve comparison between the scalar-valued and vector-valued FEM basis. The number of grid points is set to $n_{1}=101$ and $n_{2}=1048$, and the number of basis functions is $m_{1}=64$ and $m_{2}=540$.
4.5 Darcy Flow in an L-shaped domain
We investigate Darcy flow in a two-dimensional L-shaped region to validate our approach in handling multi-input scenarios. A distinctive feature of this example is that it involves a mapping from two functions to a single function, thereby highlighting the broad applicability of our method. The problem is formulated as